OpenStudy (insa):

solve by completing the square or factorization: x^2+2x+3=0

OpenStudy (anonymous):

do you know the website mathway..?

OpenStudy (perl):



OpenStudy (insa):

no

OpenStudy (anonymous):

www.mathway.com ..?

OpenStudy (insa):

no lemmee see that..

OpenStudy (insa):

hey thankxx... that website is really helpful :) :)

OpenStudy (anonymous):

glad to help.. xDDDD

OpenStudy (paki):

nice web,,, @russeldano

OpenStudy (perl):

is that website free?

OpenStudy (insa):

nope

OpenStudy (paki):

mean @insa

OpenStudy (insa):

the website ain't free :/

OpenStudy (paki):

lol... then what is the benefit of that...?

OpenStudy (neer2890):

you can check the method of completing the square here. http://www.mathsisfun.com/algebra/completing-square.html

OpenStudy (insa):

lol at least u get the correct answer

OpenStudy (anonymous):

hehe.. gee tnx.. im sharing what other people did to me.. xDDDD what did u mean free..?

OpenStudy (neer2890):

\[x ^{2}+2x+3=0\] \[x ^{2}+2x+(\frac{ 2 }{ 2 })^{2}+3-(\frac{ 2 }{ 2 })^{2}=0\] \[(x+1)^{2}+3-1=0\] \[(x+1)^{2}=-2\] \[x+1=\pm \sqrt{-2}\] \[x= \pm 2i-1\]